OpenStudy (anonymous):

Which ratio is proportional to 10/24?

HanAkoSolo (jamierox4ev3r):

basically, you are wanting to know what ratio is equivalent to this given fraction. Can this fraction be simplified? @lee_actress ?

OpenStudy (anonymous):

yes

HanAkoSolo (jamierox4ev3r):

remember, 10/24 can also be written as 10:24

OpenStudy (solomonzelman):

Hint: \[10= 2 \times 5\]\[24=2 \times 12\]

OpenStudy (solomonzelman):

@lee_actress you didn't post any choices, and you are not replying, how can we help you like this?

OpenStudy (anonymous):

Sorry. A. 35/84 B.5/6 C.5/8 D.20/36

OpenStudy (solomonzelman):

what do you think?

OpenStudy (solomonzelman):

i know the answer, but I want you to learn something.

OpenStudy (anonymous):

Could you show me how you got your answer? I have a guess.

HanAkoSolo (jamierox4ev3r):

@lee_actress , can you show us your thought process? then if you are incorrect, we can tell you where you have gone wrong

OpenStudy (anonymous):

OK

OpenStudy (solomonzelman):

Ok, I'll show you the first step. \[\huge\color{blue}{ \frac{10}{24} ~~~~->~~~~~ \frac{\color{red}{ 2 \times}5}{\color{red}{ 2 \times}12}~~~~~->~~~~ \frac{5}{12} }\]so far so good?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

I believe 5\6 is the answer because you divide 10\24 by two and you get 5\6.

OpenStudy (solomonzelman):

No it's not.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

Umm... Ok 20/36

OpenStudy (solomonzelman):

you have 5/12 try to reduce the fraction inn every option. it should be equal to 5/12. OPTIONS ARE A. 35/84 B.5/6 C.5/8 D.20/36 so which option is it?

OpenStudy (anonymous):

5/8